# A Systematic Literature Review on Detecting Software Vulnerabilities with Large Language Models

### v1

@@ 5.1.5. Dataset Programming Language
Commonly used vulnerability datasets are heavily biased towards C/C++, as shown in Table 4. This bias reflects the widespread use of these languages in technological infrastructure. Fewer datasets cover multiple programming languages, i.e., SARD, CVEfixes, CrossVul, and CleanVul; however, they are not represented in the same share. Recent studies have begun constructing datasets for less covered languages, e.g., Python (Ozturk et al., 2023) and Rust (Luo et al., 2025), contributing to broader language diversity in vulnerability research.
@@ 5.1.6. Dataset Size and Share of Vulnerable Samples
The dataset type and the labeling method used in dataset construction influence both the quality and realism, cf. Figure 4. Synthetic datasets, such as Juliet or FormAI, offer high label accuracy due to controlled injection of known vulnerability patterns but often lack the structural complexity and semantic diversity found in real-world code. In contrast, datasets derived from real-world sources provide greater realism, better reflecting practical vulnerability scenarios. However, they often suffer from label noise due to assumptions inherent in VFC-based labeling. These datasets also vary in the proportion of vulnerable samples. Highly imbalanced datasets, i.e., where vulnerabilities are underrepresented, reflect the natural rarity of vulnerabilities in real code. While realistic, such an imbalance poses challenges for training models, which may become biased towards predicting the majority class. To mitigate this challenge, preprocessing techniques are often required to ensure effective training, e.g., resampling: oversampling the minority (vulnerable) class or undersampling the majority (non-vulnerable) class to achieve a more balanced class distribution. Figure 4 visualizes this trade-off between labeling accuracy and realism, providing a practical framework for selecting datasets for different research objectives.
Matrix symbolizing the datasets’ realistic nature (type vs. vulnerability diversity and balance).
@@ 5.1.7. Dataset Multi-Class Labeling
Vulnerability detection is often framed as a binary classification task, i.e., distinguishing between vulnerable and non-vulnerable code. The ReVeal dataset provides security-related patches collected from bug tracking systems, offering paired vulnerable and fixed code samples labeled in binary form.
Many datasets go beyond binary classification by providing labels for specific vulnerability types, enabling multi-class classification. CVEfixes is a comprehensive and automatically curated dataset collected from CVE entries in the NVD. It identifies linked open-source projects and extracts function-level vulnerable and fixed code samples from VFCs. The latest release at the time of writing (v1.0.8) includes all CVEs published up to July 23, 2024, covering 272 distinct CWEs. The PairVul dataset (Du et al., 2025) takes a more controlled approach by focusing on pairs of vulnerable and non-vulnerable code with high lexical similarity. These pairs are extracted from Linux kernel CVEs, labeled using VFCs, and filtered to ensure that the associated fixes are not reverted or modified in subsequent commits. While the dataset offers high-quality and balanced samples, its scope is currently limited to selected Linux kernel vulnerabilities.
A common limitation across vulnerability datasets is the lack of coverage of recently disclosed vulnerabilities. Static datasets may quickly become outdated, reducing their relevance for real-time or forward-looking research. This highlights the need for continuously maintained and automatically extensible datasets, such as CVEfixes, which is supported by an automated pipeline.
Beyond classification, datasets such as Big-Vul, CrossVul, CVEfixes, and ReposVul offer rich metadata, including CWE descriptions and severity scores. Such metadata supports more advanced tasks such as severity estimation, root-cause reasoning, and context-aware vulnerability analysis, broadening their use beyond binary classification.
Skewed long-tail distribution of CWEs in multi-class datasets: A small number of CWE types are heavily represented, while the majority of CWE types appear only rarely.
@@ 5.2. CWE Coverage and Diversity
As discussed in prior works (Deng et al., 2024; Jiang et al., 2025; Ni et al., 2024b), real-world vulnerability datasets exhibit a characteristic long-tail distribution: A small number of CWE types are heavily represented, while the majority of CWE types appear only rarely, often with just one or two examples. This characteristic is visualized in Figure 5, which shows the count of CWEs for the 15 most and least represented CWE classes for selected datasets. For example, while CVEfixes covers 272 distinct CWE types, the majority are sparsely represented. A similar pattern is observed for DiverseVul, which covers 150 CWEs. Such skewed distributions lead to biased learning, where models disproportionately focus on a small subset of vulnerability types. MegaVul and SecurityEval represent efforts to encourage a more balanced representation. SecurityEval, in particular, provides one to six manually curated examples for 69 CWEs. These efforts represent a step towards reducing long-tail effects and developing more generalizable vulnerability detection models.
@@
Relative CWE share of each top 25 CWE within the selected dataset’s total labeled vulnerable samples.
Relative CWE share of all CWEs grouped by pillar within the selected dataset’s total labeled vulnerable samples.
The most widely represented CWEs across datasets are those associated with the pillar Improper Control of a Resource Through its Lifetime, which includes memory and resource management vulnerabilities that are commonly found in C/C++, the most dominant language across the studied datasets. This pillar is also the largest in terms of CWE count (cf. Table 5), followed by Improper Neutralization, encompassing issues related to input validation and sanitization. Other pillars are underrepresented in comparison to their relative size in the CWE view. These pillars include Improper Adherence to Coding Standards and Improper Access Control. Notably, some pillars, such as Improper Interaction Between Multiple Correctly-Behaving Entities, Protection Mechanism Failure, and Incorrect Comparison, are not represented at all among the top 25 CWEs, and are less represented in the overall CWEs covered, cf. Figure 7. These categories are inherently more difficult to capture in real-world datasets, particularly at the function level, as they often involve complex inter-component dependencies, system-level interactions, or contextual analysis that is not easily observable from isolated code snippets.
@@ 5.3. Trends and Use Cases in Dataset Utilization
Heatmap showing the use of popular datasets in studies from Jan 2020 to Mar 2025.
Figure 8 visualizes the temporal usage and usage frequency of the selected vulnerability datasets from 2020 to mid-2025 by the surveyed studies. Red markers indicate the publication dates of the datasets. The figure shows that even today, early deep learning datasets such as Devign remain in widespread use, serving as a foundational dataset for graph-based and learning-based vulnerability detection models. The introduction of Big-Vul in September 2020 further enhanced dataset diversity and volume; Big-Vul also continues to be widely adopted. The persistent use of Devign and Big-Vul is partly due to their role in enabling comparability across studies. Classic test suites such as SARD and Juliet are also frequently used, cf. Figure 8, useful for controlled evaluations due to their synthetic nature. However, these earlier datasets often lack current CWE coverage and may reinforce outdated vulnerability mappings.
@@
Another notable trend is the increasing use of newly created or custom datasets, see row ”Custom” in Figure 8. A total of 71 studies either use or introduce custom datasets, often to address perceived limitations in existing datasets. Common motivations include correcting unrealistic data distributions, expanding project and vulnerability diversity, including more recent disclosures, or improving overall dataset quality (e.g., fixing incomplete or incorrectly merged functions) (Ni et al., 2024b). This trend is also driven by domain-specific or task-aligned needs. While such efforts encourage innovation and enable tailored evaluations, they also complicate comparability and may introduce implicit biases tied to individual evaluation design choices and construction assumptions.
Heatmap showing common combination pairs of selected datasets.
@@
The choice of suitable vulnerability datasets reflects a combination of common benchmark datasets, task alignment, and system design. In total, 158 distinct datasets are used across the surveyed studies, illustrating a high degree of fragmentation. Most studies either use or introduce custom datasets, with 126 datasets appearing only once. Only eight datasets are reused in more than ten studies, highlighting the lack of a standardized suite of benchmark datasets and consistency in dataset selection. Many studies use more than one dataset, often combining datasets to increase training data diversity or to evaluate model generalization, e.g., vulnerability types and programming languages. Figure 9 highlights the most frequently used dataset pairs across all surveyed studies. Among these, the most common combinations are ReVeal and Devign, Big-Vul and Devign, as well as Big-Vul and ReVeal. These pairings often complement each other in terms of data characteristics, such as vulnerable percentage or combining datasets with broad versus narrow CWE coverage.
Matrix showing dataset overlap (same datasets used in studies) between studies across four example use cases.
@@
The use of many different datasets and combinations, however, hampers direct comparability of evaluation results, particularly when studies apply different splits for training, validation, and testing, or pre-process the same datasets differently. Figure 10(d) offers insight into overlaps and divergences in dataset usage across specific use cases. When fine-tuning LLMs for vulnerability detection, e.g., using LoRA (Low-Rank Decomposition, 17 studies), researchers often resort to large and diverse datasets, most commonly DiverseVul, Devign, Big-Vul, and ReVeal. However, many studies make use of additional, lesser-used datasets. This practice reduces comparability and makes it difficult to replicate results or isolate model performance. RAG studies (17 studies) most commonly use Devign and Big-Vul, appearing in four studies. Big-Vul, in particular, is often used for knowledge base construction of CWE entries due to the provided metadata, e.g., CWE descriptions. Five RAG studies introduce custom datasets, which hinders comparison, especially concerning the semantics of the retrieved context. Agentic approaches (8 studies) show even greater fragmentation, with only two studies aligning on PrimeVul. GNN-based hybrid approaches most frequently rely on Devign, followed by custom datasets, ReVeal, and Big-Vul. Devign is particularly suited for GNNs due to its composite code representation, encoding program control and data dependencies as heterogeneous edges within a unified graph structure. The absence of shared benchmarks within the discussed use cases highlights the need for controlled and standardized evaluation protocols.
@@ 6. Current Limitations and the Road Ahead
Despite significant progress in leveraging LLMs for software vulnerability detection, we identified several limitations in this SLR. In the following, we outline the key limitations derived from the literature and propose actionable research directions to address them. Table 6 summarizes the discussed limitations and future research opportunities.
@@ 6.1. Limitations
| • Fails to capture complex vulnerabilities with dependencies | • Augment code context |
@@
L1 Dataset and Detection Granularity
In the surveyed studies, vulnerability detection is typically applied at the level of individual functions or small program slices. This focus on function-level granularity has been influenced by the dominance of function-level vulnerability datasets (cf. Table 3) and partly by architectural limitations, particularly the restricted input lengths of earlier LLMs (Steenhoek et al., 2025a). However, methods have been proposed to enhance context, such as the work by Chen et al. (Chen et al., 2024), and LLMs now support significantly longer input contexts, e.g., 128,000 tokens for GPT4 (OpenAI et al., 2024). The focus on function-level detection fails to capture the context of and to generalize to more complex vulnerabilities that span multiple functions or files. For example, Huynh et al. (Huynh et al., 2025) investigate how varying levels of code context influence detection performance, comparing snippets enriched with comments and docstrings against full-file inputs. The findings suggest that even with extended context windows, current models still perform poorly on distributed vulnerabilities.
@@
L2 Dataset Labeling and Quality
@@
Further, most datasets suffer from a strong class imbalance, where vulnerable samples are heavily outnumbered by non-vulnerable ones (cf. Table 4), and from a long-tail distribution of vulnerability types, with few common CWEs dominating and many others being severely underrepresented (cf. Figure 5 and Figure 6). These imbalances pose significant challenges for both training balance and evaluation reliability, resulting in models achieving substantially higher performance when identifying frequently occurring vulnerabilities compared to less common ones (Huynh et al., 2025).
L3 Study Evaluation and Comparability
The lack of consistency and transparency in dataset use and evaluation protocols complicates meaningful comparison across studies. Even within the same use case, studies are often not comparable due to the use of different datasets (cf. Figure 10(d)). Further, most datasets do not provide predefined train/validate/test splits, which are essential for ensuring a consistent and fair evaluation (Shereen et al., 2024). Only a few studied datasets offer standardized splits, e.g., Draper (Russell et al., 2018), CodeXGLUE (Lu et al., 2021) (a subset of Devign (Zhou et al., 2019)), Big-Vul (Fan et al., 2020), and PrimeVul (Ding et al., 2024a). Without fixed splits, results can vary significantly, even when using the same dataset. For example, if subsequent studies fine-tune on the same dataset but apply different splits, direct performance comparisons become unreliable. In addition, several studies further modify or refine existing datasets (Zhang et al., 2024d; Yin et al., 2024; Hort et al., 2025; Gonçalves et al., 2025; Wang et al., 2024b), or augment them with custom data (Qi et al., 2024). While such adaptations are often motivated by practical needs, e.g., improving data quality or evaluating generalization, they introduce additional variability and complicate comparability with prior and future works.
L4 Up-to-Date Vulnerability Knowledge

### v2

@@ 5.1.5. Dataset Programming Language
Commonly used vulnerability datasets are heavily biased towards C/C++, as shown in Table 5. While this bias reflects the prevalence of these languages in technological infrastructure, it creates a significant imbalance, as vulnerability patterns are intrinsically linked to the programming language. Qian et al. (Qian et al., 2025) highlight that unmanaged languages such as C/C++ are prone to memory safety violations (e.g., buffer overflows), whereas managed languages such as JavaScript or Python frequently suffer from logic bugs. In a comparative analysis of interpreted languages (Java, JavaScript, Python, and Ruby), Ruohonen (Ruohonen, 2021) observes that while certain weaknesses such as cross-site scripting and input validation errors are ubiquitous across all ecosystems, language-specific trends emerge. For instance, cryptographic issues were found to be prevalent in JavaScript but significantly less common in Java or Python (Ruohonen, 2021). Consequently, the predominance of C/C++ risks over-optimizing models for memory safety violations while potentially neglecting the distinct vulnerability landscapes of modern web languages.
Fewer datasets cover multiple programming languages, i.e., SARD, CVEfixes, CrossVul, and CleanVul; however, they are not represented in the same share. Recent studies have begun constructing datasets for less covered languages, e.g., Python (Ozturk et al., 2023) and Rust (Luo et al., 2025), contributing to broader language diversity in vulnerability research.
@@ 5.1.6. Dataset Size and Share of Vulnerable Samples
The dataset type and the labeling method used in dataset construction influence both the quality and realism, cf. Figure 7. Synthetic datasets, such as Juliet or FormAI, offer high label accuracy due to controlled injection of known vulnerability patterns but often lack the structural complexity and semantic diversity found in real-world code. In contrast, datasets derived from real-world sources provide greater realism, better reflecting practical vulnerability scenarios. However, they often suffer from label noise due to assumptions inherent in VFC-based labeling. These datasets also vary in the proportion of vulnerable samples. Highly imbalanced datasets, i.e., where vulnerabilities are underrepresented, reflect the natural rarity of vulnerabilities in real code. While realistic, such an imbalance poses challenges for training models, which may become biased towards predicting the majority class. To mitigate this challenge, preprocessing techniques are often required, e.g., resampling: oversampling the minority (vulnerable) class or undersampling the majority (non-vulnerable) class to achieve a more balanced class distribution. Figure 7 visualizes this trade-off between labeling accuracy and realism, providing a practical framework for selecting datasets for different research objectives.
@@ 5.1.7. Dataset Multi-Class Labeling
Vulnerability detection is often framed as a binary classification task, i.e., distinguishing between vulnerable and non-vulnerable code. For example, the ReVeal dataset provides security-related patches collected from bug tracking systems, offering paired vulnerable and fixed code samples labeled in binary form.
Many datasets go beyond binary classification by providing labels for specific vulnerability types, enabling multi-class and multi-label classification. CVEfixes is a comprehensive and automatically curated dataset collected from CVE entries in the NVD. It identifies linked open-source projects and extracts function-level vulnerable and fixed code samples from VFCs. The latest release at the time of writing (v1.0.8) includes all CVEs published up to July 23, 2024, covering 272 distinct CWEs. The PairVul dataset (Du et al., 2025a) takes a more controlled approach by focusing on pairs of vulnerable and non-vulnerable code with high lexical similarity. These pairs are extracted from Linux kernel CVEs, labeled using VFCs, and filtered to ensure that the associated fixes are not reverted or modified in subsequent commits. While the dataset offers high-quality and balanced samples, its scope is currently limited to selected Linux kernel vulnerabilities.
A common limitation across vulnerability datasets is the lack of coverage of recently disclosed vulnerabilities. Static datasets may quickly become outdated, reducing their relevance for real-time or forward-looking research. In addition, this temporal lag introduces a critical validity threat for LLMs in the form of data leakage. Since widely used static datasets are likely contained within the pre-training corpora of LLMs and new models are continuously emerging, high performance on these benchmarks may indicate memorization rather than genuine generalization. This challenge highlights the need for continuously maintained and extensible datasets, e.g., CVEfixes, which is supported by an automated pipeline, to ensure evaluation against novel, unseen vulnerabilities.
Datasets such as Big-Vul, CrossVul, CVEfixes, and ReposVul further offer rich metadata, including CWE descriptions and severity scores. Such metadata supports diverse generative tasks, for example, root-cause reasoning and context-aware vulnerability analysis, broadening their use beyond classification.
@@ 5.2. CWE Coverage and Diversity
As discussed in prior works (Deng et al., 2024; Jiang et al., 2025b; Ni et al., 2024), real-world vulnerability datasets exhibit a characteristic long-tail distribution: A small number of CWE types are heavily represented, while the majority of CWE types appear only rarely, often with just one or two examples. This characteristic is visualized in Figure 8, which shows the count of CWEs for the 15 most and least represented CWE classes for selected datasets. For example, while CVEfixes covers 272 distinct CWE types, the majority are sparsely represented. A similar pattern is observed for DiverseVul, which covers 150 CWEs. Such skewed distributions lead to biased learning, where models disproportionately focus on a small subset of vulnerability types. MegaVul and SecurityEval represent efforts to encourage a more balanced representation. SecurityEval, in particular, provides one to six manually curated examples for 69 CWEs. These efforts represent a step towards reducing long-tail effects and developing more generalizable vulnerability detection models.
@@
The most widely represented CWEs across datasets are those associated with the pillar Improper Control of a Resource Through its Lifetime, which includes memory and resource management vulnerabilities that are commonly found in C/C++, the most dominant language across the studied datasets (cf. discussion in Section 5.1.5). This pillar is also the largest in terms of CWE count (cf. Table 6), followed by Improper Neutralization, encompassing issues related to input validation and sanitization. Other pillars are underrepresented in comparison to their relative size in the CWE view. These pillars include Improper Adherence to Coding Standards and Improper Access Control. Notably, some pillars, such as Improper Interaction Between Multiple Correctly-Behaving Entities, Protection Mechanism Failure, and Incorrect Comparison, are not represented at all among the Top 25 CWEs, and are less represented in the overall CWEs covered, cf. Figure 10. These categories are inherently more difficult to capture in real-world datasets, particularly at the function level, as they often involve complex inter-component dependencies, system-level interactions, or contextual analysis that is not easily observable from isolated code snippets.
@@ 5.3. Trends and Use Cases in Dataset Utilization
Figure 11 visualizes the temporal usage and usage frequency of the selected vulnerability datasets from Jan-2020 to Nov-2025 by the surveyed studies. Red markers indicate the publication dates of the datasets. The Figure shows that even today, early deep learning datasets such as Devign, which serves as a foundational dataset for graph-based and learning-based vulnerability detection models, and Big-Vul remain in widespread use. Classic test suites such as SARD and Juliet are also frequently used, cf. Figure 11, useful for controlled evaluations due to their synthetic nature. However, these earlier datasets often lack current CWE coverage, may reinforce outdated vulnerability mappings, and may already be part of training corpora.
@@
Another notable trend is the increasing use of newly created or custom datasets, see row ”Custom” in Figure 11. A total of 87 studies either use or introduce custom datasets, often to address perceived limitations in existing datasets. Common motivations include correcting unrealistic data distributions, expanding project and vulnerability diversity, including more recent disclosures, or improving overall dataset quality (e.g., fixing incomplete or incorrectly merged functions) (Ni et al., 2024). This trend is also driven by domain-specific or task-aligned needs. While such efforts encourage innovation and enable tailored evaluations, they also complicate comparability and may introduce implicit biases tied to individual evaluation design choices and construction assumptions.
@@
The choice of suitable vulnerability datasets reflects a combination of common benchmark datasets, task alignment, and system design. However, the use of various datasets across the surveyed studies illustrates a high degree of fragmentation. Most studies either use or introduce custom datasets. Only nine datasets are reused in more than ten studies, highlighting the lack of a standardized suite of benchmark datasets and consistency in dataset selection. Many studies use more than one dataset, often combining datasets to increase training data diversity or to evaluate model generalization, e.g., vulnerability types and programming languages. Figure 12 highlights the most frequently used dataset pairs across all surveyed studies. Among these, the most common combinations are Devign and ReVeal, Devign and Big-Vul, as well as Big-Vul and ReVeal. These pairings often complement each other in terms of data characteristics, such as vulnerable percentage or combining datasets with broad versus narrow CWE coverage.
@@
The use of many different datasets and combinations, however, hampers direct comparability of evaluation results, particularly when studies apply different splits for training, validation, and testing, or pre-process the same datasets differently. Figure 13 offers insight into overlaps and divergences in dataset usage across specific use cases. When fine-tuning LLMs for vulnerability detection, e.g., using LoRA (Low-Rank Decomposition, 27 studies), researchers often resort to large and diverse datasets, most commonly DiverseVul, Big-Vul, PrimeVul, and Devign. However, many studies make use of additional, lesser-used datasets or introduce custom datasets (ten studies). This practice reduces comparability and makes it difficult to replicate results or isolate model performance. RAG studies (25 studies) most commonly use Big-Vul and Devign, appearing in six and five studies, respectively. Big-Vul, in particular, is often used for knowledge base construction of CWE entries due to the provided metadata, e.g., CWE descriptions. Eight RAG studies introduce custom datasets, which hinder comparison, especially concerning the semantics of the retrieved context. Agentic approaches (10 studies) show even greater fragmentation, with only four studies aligning on PrimeVul. GNN-based hybrid approaches (20 studies) most frequently rely on Devign, followed by custom datasets, Big-Vul, and ReVeal. Devign is particularly suited for GNNs due to its composite code representation, encoding program control and data dependencies as heterogeneous edges within a unified graph structure. The absence of shared benchmarks within the discussed use cases highlights the need for controlled and standardized evaluation protocols.
@@ 6. Current Limitations and the Road Ahead
Despite progress in leveraging LLMs for software vulnerability detection, we identified several remaining limitations in this SLR. In the following, we outline the key limitations derived from the surveyed studies and propose actionable research directions to address them. Table 7 summarizes the discussed limitations and future research opportunities.
@@ 6.1. Limitations
| • Failure to capture complex vulnerabilities with dependencies | • Augment code context |
@@
| • Missing reasoning ground truth | • Reasoning-oriented datasets and benchmarks |
@@
| • Data leakage | • Build new datasets considering knowledge cutoff |
@@
6.1.1. L1 Dataset and Detection Granularity
In the surveyed studies, vulnerability detection is typically applied at the level of individual functions or small program slices. This focus on function-level granularity has been influenced by the dominance of function-level vulnerability datasets (cf. Table 4) and partly by architectural limitations, particularly the restricted input lengths of earlier LLMs (Steenhoek et al., 2025a). However, methods have been proposed to enhance context, such as the work by Chen et al. (Chen et al., 2024), and LLMs now support significantly longer input contexts, e.g., 128,000 tokens for GPT4 (OpenAI et al., 2024). Despite these capabilities, the continued focus on function-level datasets fails to capture the context necessary to generalize to complex vulnerabilities that span multiple functions or files. For example, Huynh et al. (Huynh et al., 2025) investigate how varying levels of code context influence detection performance, comparing snippets enriched with comments and docstrings against full-file inputs. The findings suggest that even with extended context windows, current models still perform poorly on distributed vulnerabilities.
@@ 6.1.1. L1 Dataset and Detection Granularity
6.1.2. L2 Dataset Labeling and Quality
@@
Moreover, as research shifts toward generative tasks, a critical gap of missing reasoning ground truth emerges. While existing datasets typically provide the label and possibly a description of the vulnerability, they rarely provide a verified chain of thought reasoning the root cause. This limitation makes it difficult to quantitatively benchmark the quality of LLM-generated reasoning without expensive human evaluation.
Further, most datasets exhibit a strong class imbalance, where vulnerable samples are heavily outnumbered by non-vulnerable ones (cf. Table 5), and a long-tail distribution of vulnerability types, with few CWEs dominating and many others being severely underrepresented (cf. Figure 8 and Figure 9). These imbalances pose significant challenges for both training balance and evaluation reliability, causing models to overfit to frequent vulnerabilities while failing on rare ones (Huynh et al., 2025).
Beyond these traditional labeling and quality issues, LLMs introduce the unique challenge of data leakage. Unlike models trained on closed datasets, LLMs are pre-trained on vast open-source corpora. Consequently, popular benchmarks (e.g., Devign, Big-Vul) are likely contained within the pre-training corpus, rendering standard train/test splits unreliable. This challenge is compounded by the knowledge cut-off; without continuously updated datasets that post-date the model’s training window, it is difficult to evaluate whether an LLM is detecting a vulnerability via reasoning or simply recalling historical data it has already seen.
6.1.3. L3 Study Evaluation and Comparability
The lack of consistency and transparency in dataset use and evaluation protocols complicates meaningful comparison across studies. Even within the same use case, studies are often not comparable due to the use of different datasets (cf. Figure 13). Further, most datasets do not provide predefined train/validate/test splits, which are essential for ensuring a consistent and fair evaluation (Shereen et al., 2024). Only a few studied datasets offer standardized splits, e.g., Draper (Russell et al., 2018), CodeXGLUE (Lu et al., 2021) (a subset of Devign (Zhou et al., 2019)), Big-Vul (Fan et al., 2020), and PrimeVul (Ding et al., 2024a). Without fixed splits, results can vary significantly, even when using the same dataset. For example, if subsequent studies fine-tune on the same dataset but apply different splits, direct performance comparisons become unreliable. In addition, several studies further modify or refine existing datasets (Zhang et al., 2024c; Yin et al., 2024; Hort et al., 2025; Gonçalves et al., 2025; Wang et al., 2025c), or augment them with custom data (Qi et al., 2024). While such adaptations are often motivated by practical needs, e.g., improving data quality or evaluating generalization, they introduce additional variability and complicate comparability with prior and future works.
6.1.4. L4 Up-to-Date Vulnerability Knowledge